Jefferson Lab HALOG ARCHIVES Logbook for June 05 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 147668 06/30/05 23:58 apar Start_Parity_Run_3475__CODA_Config=Beamline
  • 147667 06/30/05 23:57 apar End_of_Parity_Run_3474__CODA_Config=Beamline
  • 147666 06/30/05 23:55 R.A. Lindgre shift summary
  • 147665 06/30/05 23:50 apar Start_Parity_Run_3474__CODA_Config=Beamline
  • 147664 06/30/05 23:49 apar End_of_Parity_Run_3473__CODA_Config=Beamline
  • 147663 06/30/05 23:41 apar Start_Parity_Run_3473__CODA_Config=Beamline
  • 147662 06/30/05 23:40 apar End_of_Parity_Run_3472__CODA_Config=Beamline
  • 147661 06/30/05 23:38 apar Start_Parity_Run_3472__CODA_Config=Beamline
  • 147660 06/30/05 23:37 apar End_of_Parity_Run_3471__CODA_Config=Beamline
  • 147659 06/30/05 23:27 apar Start_Parity_Run_3471__CODA_Config=Beamline
  • 147658 06/30/05 23:26 apar End_of_Parity_Run_3470__CODA_Config=Beamline
  • 147656 06/30/05 23:19 apar Start_Parity_Run_3470__CODA_Config=Beamline
  • 147655 06/30/05 23:18 apar End_of_Parity_Run_3469__CODA_Config=Beamline
  • 147653 06/30/05 23:16 apar Start_Parity_Run_3469__CODA_Config=Beamline
  • 147652 06/30/05 22:50 Armando Acha target
  • 147651 06/30/05 22:49 R.A. LIndgre Waterfall target stuck at home position and magnet problems
  • 147622 06/30/05 18:46 apar End_of_Parity_Run_3468__CODA_Config=Beamline
  • 147620 06/30/05 18:32 apar Start_Parity_Run_3468__CODA_Config=Beamline
  • 147619 06/30/05 18:31 apar End_of_Parity_Run_3467__CODA_Config=Beamline
  • 147615 06/30/05 17:51 apar Start_Parity_Run_3467__CODA_Config=Beamline
  • 147614 06/30/05 17:50 apar End_of_Parity_Run_3466__CODA_Config=Beamline
  • 147612 06/30/05 17:45 apar Start_Parity_Run_3466__CODA_Config=Beamline
  • 147604 06/30/05 17:25 apar End_of_Parity_Run_3465__CODA_Config=Beamline
  • 147602 06/30/05 17:22 apar Start_Parity_Run_3465__CODA_Config=Beamline
  • 147596 06/30/05 17:04 apar End_of_Parity_Run_3464__CODA_Config=Beamline
  • 147595 06/30/05 17:02 apar Start_Parity_Run_3464__CODA_Config=Beamline
  • 147594 06/30/05 16:53 spiegel Hall A Tech's
  • 147593 06/30/05 16:50 apar End_of_Parity_Run_3463__CODA_Config=Beamline
  • 147592 06/30/05 16:47 apar Start_Parity_Run_3463__CODA_Config=Beamline
  • 147591 06/30/05 16:39 apar End_of_Parity_Run_3462__CODA_Config=Beamline
  • 147590 06/30/05 16:36 apar Start_Parity_Run_3462__CODA_Config=Beamline
  • 147589 06/30/05 16:16 R.A. Lindgre Polarity change underway
  • 147587 06/30/05 16:01 B. Craver Day shift summary
  • 147586 06/30/05 15:49 apar End_of_Parity_Run_3461__CODA_Config=happex2
  • 147585 06/30/05 14:54 apar Start_Parity_Run_3461__CODA_Config=happex2
  • 147583 06/30/05 14:50 apar End_of_Parity_Run_3460__CODA_Config=happex2
  • 147582 06/30/05 14:21 LeRose June 30 Swing Shift Plan
  • 147579 06/30/05 13:55 apar Start_Parity_Run_3460__CODA_Config=happex2
  • 147578 06/30/05 13:46 B. Craver Hall A tools and BCM stripchart
  • 147577 06/30/05 13:42 B. Craver SLI and accelerator screen shot
  • 147568 06/30/05 10:43 Adaq SLI screen shot
  • 147565 06/30/05 09:59 apar End_of_Parity_Run_3459__CODA_Config=happex2_min
  • 147564 06/30/05 09:51 apar Start_Parity_Run_3459__CODA_Config=happex2_min
  • 147559 06/30/05 08:09 E. Cisbani Owl Shift Summary 06/30
  • 147558 06/30/05 08:09 E. Cisbani 8hrs BCM stripchart
  • 147552 06/30/05 06:45 E. Cisbani SLI screen + Accelerator parameters
  • 147551 06/30/05 06:11 E. Cisbani Hall A tools screen
  • 147541 06/30/05 03:08 E. Cisbani SLI screen
  • 147535 06/30/05 01:16 E. Cisbani Energy fluctuation due to fast feedback standby
  • 147531 06/30/05 00:04 S. Sirca Evening shift summary
  • 147530 06/30/05 00:02 S. Sirca SLI fixed
  • Wednesday

  • 147528 06/29/05 23:56 S. Sirca 8-hour BCM etc screen
  • 147527 06/29/05 23:42 S. Sirca Lost SLI information
  • 147521 06/29/05 21:23 S. Sirca Beam restored
  • 147519 06/29/05 21:05 S. Sirca Beam off for injector work
  • 147517 06/29/05 20:55 S. Sirca SLI snapshot - picture
  • 147516 06/29/05 20:54 S. Sirca SLI snapshot before injector work
  • 147508 06/29/05 18:32 S. Sirca Beam restored
  • 147507 06/29/05 18:29 S. Sirca Electron gun lost
  • 147503 06/29/05 17:40 S. Sirca General Tools snapshot
  • 147502 06/29/05 17:38 S. Sirca SLI snapshot
  • 147497 06/29/05 16:04 C. C. Chang Shift summary
  • 147495 06/29/05 15:42 apar End_of_Parity_Run_3458__CODA_Config=Beamline
  • 147494 06/29/05 15:28 C. C. Chang Beam energy changed
  • 147490 06/29/05 15:17 C. C. Chang SLI
  • 147489 06/29/05 15:13 C. C. Chang Right septum is back
  • 147488 06/29/05 15:08 apar Start_Parity_Run_3458__CODA_Config=Beamline
  • 147487 06/29/05 14:35 C. C. Chang Right septum magnet trip
  • 147485 06/29/05 14:24 C. C. Chang Magnet strip charts
  • 147482 06/29/05 14:00 C. C. Chang BCM strip charts
  • 147481 06/29/05 13:59 C. C. Chang Hall A Tools
  • 147476 06/29/05 12:03 C. C. Chang SLI
  • 147475 06/29/05 12:00 C. C. Chang Energy spread
  • 147474 06/29/05 11:41 C. C. Chang SLI
  • 147473 06/29/05 11:39 C. C. Chang Beam back
  • 147471 06/29/05 11:12 C. C. Chang RF separator problem, no beam
  • 147465 06/29/05 10:31 C. C. Chang Energy spread
  • 147462 06/29/05 09:47 apar End_of_Parity_Run_3457__CODA_Config=happex2_min
  • 147461 06/29/05 09:42 apar Start_Parity_Run_3457__CODA_Config=happex2_min
  • 147460 06/29/05 09:41 apar End_of_Parity_Run_3456__CODA_Config=happex2_min
  • 147459 06/29/05 09:41 apar Start_Parity_Run_3456__CODA_Config=happex2_min
  • 147453 06/29/05 08:01 E. Cisbani Owl Shift Summary 06/29
  • 147449 06/29/05 07:09 E. Cisbani Fast Feedback in standby, restarted
  • 147446 06/29/05 06:21 E. Cisbani Hall A Tools screen
  • 147445 06/29/05 06:20 E. Cisbani SLI screen
  • 147444 06/29/05 06:19 E. Cisbani Energy fluctuation fixed
  • 147443 06/29/05 06:06 E. Cisbani Large Beam Energy fluctuation
  • 147436 06/29/05 04:01 Armando Acha beam dots
  • 147433 06/29/05 02:48 E. Cisbani Beam back, energy spread fine
  • 147432 06/29/05 02:43 E. Cisbani SLI screen
  • 147430 06/29/05 02:19 E. Cisbani Beam tuning
  • 147428 06/29/05 02:08 E. Cisbani Bad energy spread
  • 147426 06/29/05 02:04 Armando Acha raster
  • 147423 06/29/05 01:52 E. Cisbani Beam back
  • 147422 06/29/05 01:35 E. Cisbani Beam off due to power loss in W4
  • 147421 06/29/05 00:21 E. Cisbani No beam RCS/IOC lost
  • 147418 06/29/05 00:02 LeRose Shift summary
  • Tuesday

  • 147409 06/28/05 21:08 LeRose Hall A Tools
  • 147408 06/28/05 21:07 LeRose SLI
  • 147407 06/28/05 21:04 LeRose Nice Beam
  • 147402 06/28/05 18:56 LeRose Right Septum back
  • 147400 06/28/05 18:27 LeRose Right septum Trip
  • 147396 06/28/05 17:46 LeRose SLI
  • 147395 06/28/05 17:43 LeRose Espread
  • 147394 06/28/05 17:34 LeRose Espread
  • 147391 06/28/05 16:36 LeRose Beam Back
  • 147389 06/28/05 16:23 LeRose No Beam/RF seperator trip
  • 147385 06/28/05 16:02 reitz shift summary day shift
  • 147377 06/28/05 13:16 Reitz eye of the bull
  • 147332 06/28/05 10:00 reitz SLI screen shot
  • 147328 06/28/05 09:04 M.Iodice Right Septum Trip
  • 147322 06/28/05 08:00 E. Cisbani Owl Shift summary
  • 147321 06/28/05 06:40 E. Cisbani RF cavity problem fixed
  • 147320 06/28/05 06:01 E. Cisbani Beam off, RF failure under investigation
  • 147319 06/28/05 04:39 Higinbotham Position Fast Feed Back
  • 147318 06/28/05 04:31 E. Cisbani Beam off
  • 147316 06/28/05 03:56 E. Cisbani SLI screen
  • 147313 06/28/05 03:33 E. Cisbani Waterfall target speed drift
  • 147312 06/28/05 03:02 Armando Acha beam
  • 147309 06/28/05 02:47 Armando Acha beam
  • 147308 06/28/05 02:27 E. Cisbani BCM spread, large fluctuation reduced
  • 147305 06/28/05 02:15 E. Cisbani BCM spread
  • 147304 06/28/05 01:43 E. Cisbani SLI screen
  • 147299 06/28/05 00:19 E. Cisbani SLI & Accelerator Parameters screens
  • 147298 06/28/05 00:01 E.Chudakov Shift summary
  • Monday

  • 147296 06/27/05 23:30 E.Chudakov Beam is back
  • 147294 06/27/05 21:49 E.Chudakov Beam stopped in order to tune Hall C
  • 147290 06/27/05 21:29 E.Chudakov Energy spread
  • 147289 06/27/05 20:47 E.Chudakov Resumed production with the Moller quads off
  • 147283 06/27/05 20:27 folts tech/beam line vacuum
  • 147282 06/27/05 20:19 E.Chudakov Spot size with Moller quads off
  • 147281 06/27/05 19:59 E.Chudakov Wrong beam optics
  • 147276 06/27/05 19:20 E.Chudakov SLI energy spread
  • 147275 06/27/05 19:15 E.Chudakov CW beam in the hall
  • 147272 06/27/05 18:31 E.Chudakov Going to beam permit
  • 147270 06/27/05 15:56 reitz shift summary (day)
  • 147266 06/27/05 09:24 Markowitz magnets set
  • 147265 06/27/05 07:36 Markowitz left arm magnets / iocha14
  • Sunday

  • 147264 06/26/05 20:02 folts tech
  • 147263 06/26/05 18:30 folts tech
  • 147262 06/26/05 17:18 cusanno updated missing energy spectrum
  • 147261 06/26/05 15:37 Markowitz target update / shifts cancelled tonight
  • 147260 06/26/05 14:06 Markowitz target update / shift schedule
  • 147257 06/26/05 13:08 Markowitz target leak
  • 147254 06/26/05 08:03 E. Cisbani Owl Shift summary
  • 147253 06/26/05 06:49 Markowitz Vacuum expert from accelerator coming in
  • 147252 06/26/05 05:09 E. Cisbani Access required for vacuum fix
  • 147251 06/26/05 04:39 E. Cisbani Beam off due to vacuum valves
  • 147249 06/26/05 04:21 E. Cisbani SLI screen
  • 147244 06/26/05 02:52 E. Cisbani SLI screen
  • 147240 06/26/05 01:54 E. Cisbani Beam Back
  • 147239 06/26/05 00:04 LeRose Shift Summary
  • Saturday

  • 147237 06/25/05 22:24 LeRose Beam Off
  • 147234 06/25/05 22:18 LeRose Energy Spread Back
  • 147232 06/25/05 21:49 LeRose Energy spread
  • 147228 06/25/05 21:29 LeRose BCM/spreads/septum
  • 147227 06/25/05 21:26 LeRose Hall A Tools
  • 147226 06/25/05 21:25 LeRose SLI
  • 147225 06/25/05 21:21 LeRose HALLA:p bad
  • 147222 06/25/05 20:12 LeRose Beam to Hall C
  • 147220 06/25/05 19:45 PPamela DAQ ROC7 reboot
  • 147218 06/25/05 19:14 LeRose 55 microamps
  • 147215 06/25/05 18:54 LeRose BCM/spread
  • 147214 06/25/05 18:51 LeRose SLI
  • 147211 06/25/05 17:36 LeRose 50 microamps
  • 147210 06/25/05 17:32 LeRose Septum Strip Chart
  • 147208 06/25/05 17:30 LeRose Septum Back
  • 147207 06/25/05 16:55 LeRose Right Septum Trip
  • 147201 06/25/05 16:05 bogdanw shift summary
  • 147194 06/25/05 13:31 bogdanw right septa
  • 147188 06/25/05 11:45 bogdanw SLI
  • 147181 06/25/05 08:41 bogdanw SLI
  • 147180 06/25/05 08:40 bogdanw septa screens
  • 147179 06/25/05 08:38 bogdanw Hall A tools
  • 147178 06/25/05 08:36 bogdanw RICH screen
  • 147176 06/25/05 08:13 M.Iodice Reduced Waterfall pump speed
  • 147175 06/25/05 08:05 E. Cisbani Owl Shift Summary 06/25
  • 147172 06/25/05 07:35 E. Cisbani SLI screen
  • 147169 06/25/05 06:55 E. Cisbani SLI screen
  • 147168 06/25/05 06:54 E. Cisbani RICH snapshot
  • 147164 06/25/05 06:26 E. Cisbani No End of Run 4890 e-log entry
  • 147161 06/25/05 06:04 E. Cisbani Right Septum Chart and Target Tickness
  • 147160 06/25/05 05:50 E. Cisbani SLI screen
  • 147159 06/25/05 05:36 E. Cisbani Hall A Current Monitor
  • 147155 06/25/05 04:33 E. Cisbani SLI screen
  • 147154 06/25/05 04:20 E. Cisbani Target Thickness Changed
  • 147153 06/25/05 04:16 E. Cisbani ROC16 rebooted and CODA restarted
  • 147148 06/25/05 03:31 E. Cisbani SLI screen
  • 147145 06/25/05 02:23 E. Cisbani SLI screen
  • 147142 06/25/05 01:36 E. Cisbani ROC12 rebooted and CODA restarted
  • 147139 06/25/05 00:46 E. Cisbani SLI screen
  • 147136 06/25/05 00:06 S. Marrone Shift Summary
  • Friday

  • 147126 06/24/05 20:33 S. Marrone SLI
  • 147125 06/24/05 20:31 S. Marrone Strip Chart
  • 147115 06/24/05 16:30 S. Marrone Shift Summary
  • 147110 06/24/05 14:38 apar End_of_Parity_Run_3455__CODA_Config=happex2_min
  • 147109 06/24/05 14:33 apar Start_Parity_Run_3455__CODA_Config=happex2_min
  • 147106 06/24/05 14:29 R. Michaels EPICS error messages
  • 147105 06/24/05 14:11 R. Michaels getscaler change
  • 147102 06/24/05 13:41 apar End_of_Parity_Run_3454__CODA_Config=happex2_min
  • 147101 06/24/05 13:33 apar Start_Parity_Run_3454__CODA_Config=happex2_min
  • 147100 06/24/05 13:30 cusanno beam back
  • 147099 06/24/05 13:25 cusanno beam off
  • 147098 06/24/05 13:13 apar End_of_Parity_Run_3453__CODA_Config=happex2_min
  • 147097 06/24/05 13:07 apar Start_Parity_Run_3453__CODA_Config=happex2_min
  • 147096 06/24/05 12:57 cusanno Hall A tools screenshot
  • 147095 06/24/05 12:53 cusanno energy lock ok, beam back
  • 147093 06/24/05 11:37 cusanno energy lock problems, beam off
  • 147091 06/24/05 10:45 cusanno beam back
  • 147089 06/24/05 10:43 cusanno No beam
  • 147087 06/24/05 10:17 cusanno Right septum screenshot
  • 147086 06/24/05 10:13 cusanno BCM/SLI screenshot
  • 147085 06/24/05 10:12 cusanno SLI screenshot
  • 147080 06/24/05 09:04 cusanno BCM/SLI screen shot
  • 147079 06/24/05 09:02 cusanno SLI screenshot
  • 147075 06/24/05 08:10 E. Cisbani Owl Shift Summary
  • 147072 06/24/05 07:45 E. Cisbani O2 fluctuation in RICH
  • 147071 06/24/05 07:11 E. Cisbani SLI screen
  • 147066 06/24/05 06:09 E. Cisbani SLI screen + GAP
  • 147065 06/24/05 06:06 E. Cisbani Hall A Tools screen
  • 147062 06/24/05 05:06 E. Cisbani SLI screen
  • 147061 06/24/05 05:03 E. Cisbani Right Septum Magnet Tripped
  • 147056 06/24/05 03:29 E. Cisbani SLI screen
  • 147055 06/24/05 03:28 E. Cisbani Hall A Current Monitor
  • 147054 06/24/05 03:11 E. Cisbani ROC16 rebooted, CODA restarted
  • 147051 06/24/05 02:27 E. Cisbani SLI screen
  • 147050 06/24/05 02:23 E. Cisbani ROC12 rebooted, CODA restarted
  • 147045 06/24/05 01:13 E. Cisbani SLI screen
  • 147042 06/24/05 00:12 E. Cisbani SLI screen
  • 147041 06/24/05 00:01 Holmstrom Swing Shift Summary
  • Thursday

  • 147039 06/23/05 23:58 Aniol det and phys replays done
  • 147036 06/23/05 23:04 Holmstrom SLI screen
  • 147031 06/23/05 22:00 LeRose Owl Shift Instructions
  • 147030 06/23/05 21:40 Holmstrom SLI Screen
  • 147024 06/23/05 20:01 Aniol run 4843,4844 ok
  • 147021 06/23/05 19:13 aniol RICH screen
  • 147018 06/23/05 18:29 Holmstrom SLI Strip Tool
  • 147016 06/23/05 17:52 LeRose Correction
  • 147015 06/23/05 16:11 LeRose Shift Summary
  • 147014 06/23/05 16:02 apar End_of_Parity_Run_3452__CODA_Config=happex2_min
  • 147013 06/23/05 16:01 apar Start_Parity_Run_3452__CODA_Config=happex2_min
  • 147012 06/23/05 15:49 E.Chudakov Control access
  • 147011 06/23/05 15:04 LeRose Left Arm Back
  • 147010 06/23/05 14:26 LeRose Left Arm Controls Frozen
  • 147009 06/23/05 12:51 LeRose Magnets back
  • 147008 06/23/05 12:22 LeRose Magnets Tripped
  • 147007 06/23/05 11:34 Fansler/Spie Techs
  • 147004 06/23/05 07:59 Raue shift summary
  • 147000 06/23/05 07:21 Armando Acha RICH parameters
  • 146999 06/23/05 07:15 Armando Acha RICH strips
  • 146998 06/23/05 06:58 Raue Beam energy spread
  • 146996 06/23/05 06:14 Raue bad beam energy spread
  • 146995 06/23/05 05:58 Armando Acha beam spot raster
  • 146994 06/23/05 05:25 Raue right septum tripped
  • 146993 06/23/05 04:12 Raue waiting
  • 146992 06/23/05 03:08 Raue still no beam
  • 146991 06/23/05 02:18 PPamela Replay, not-so-good beam, BPM y vs x
  • 146981 06/23/05 00:02 Holmstrom Swing Shift Summary
  • Wednesday

  • 146980 06/22/05 22:46 Holmstrom RICH Strip Chart
  • 146976 06/22/05 20:34 apar End_of_Parity_Run_3451__CODA_Config=CountHouse
  • 146975 06/22/05 20:28 apar Start_Parity_Run_3451__CODA_Config=CountHouse
  • 146974 06/22/05 20:27 apar End_of_Parity_Run_3450__CODA_Config=CountHouse
  • 146973 06/22/05 20:24 apar Start_Parity_Run_3450__CODA_Config=CountHouse
  • 146971 06/22/05 20:19 Holmstrom CHL Crash
  • 146967 06/22/05 19:21 Holmstrom BCM Spread Strip Chart
  • 146966 06/22/05 19:16 Aniol RICH screen
  • 146962 06/22/05 18:07 apar End_of_Parity_Run_3449__CODA_Config=Lumi
  • 146961 06/22/05 18:04 apar Start_Parity_Run_3449__CODA_Config=Lumi
  • 146960 06/22/05 17:33 Holmstrom SLI Snapshot
  • 146956 06/22/05 16:07 C. C. Chang Shift summary
  • 146953 06/22/05 15:57 C. C. Chang Right septum is back
  • 146951 06/22/05 15:23 C. C. Chang Right septum trip
  • 146947 06/22/05 14:18 R. Michaels High DT, stopped run 4825
  • 146945 06/22/05 14:14 R. Michaels ROC7 errors, changed
  • 146944 06/22/05 14:12 C. C. Chang SLI screen shot
  • 146943 06/22/05 14:08 C. C. Chang Septa power supplies screen shots
  • 146942 06/22/05 14:05 C. C. Chang RICH strip charts
  • 146941 06/22/05 14:04 C. C. Chang Right septum strip charts
  • 146940 06/22/05 14:02 C. C. Chang Magnet strip charts
  • 146939 06/22/05 13:59 C. C. Chang Hall A Tools
  • 146937 06/22/05 13:35 C. C. Chang Alarm handler
  • 146932 06/22/05 12:19 R. Michaels SLI screen
  • 146931 06/22/05 11:58 R. Michaels scaler checkout
  • 146928 06/22/05 11:05 C. C. Chang Beam back
  • 146923 06/22/05 09:50 C. C. Chang No beam
  • 146922 06/22/05 09:30 C. C. Chang All magnets are back
  • 146921 06/22/05 09:30 ewing techs
  • 146920 06/22/05 08:59 C. C. Chang Right septum trip
  • 146919 06/22/05 08:55 C. C. Chang Hall access
  • 146918 06/22/05 08:48 C. C. chang Magnets trip
  • 146916 06/22/05 08:45 R. Michaels power cycle ROC3 fixes it
  • 146913 06/22/05 07:55 R. Michaels ROC3 slot 9 not responding
  • 146907 06/22/05 07:46 Raue shift summary
  • 146901 06/22/05 07:06 Raue dead time problem
  • 146886 06/22/05 06:05 Raue SLI Screen
  • 146882 06/22/05 05:54 Armando Acha dead time
  • 146881 06/22/05 05:40 Armando Acha Missing det histograms for L arm
  • 146880 06/22/05 04:45 Raue SLI screen
  • 146879 06/22/05 04:45 Raue dead time
  • 146876 06/22/05 04:34 Armando Acha rich rings
  • 146875 06/22/05 03:49 PPamela Replay trouble continued
  • 146873 06/22/05 03:42 Raue alarm problem?
  • 146872 06/22/05 03:19 Raue SLI screen
  • 146871 06/22/05 03:09 PPamela Replay trouble
  • 146868 06/22/05 02:02 Markowitz replay from adaql1 only -- something weird
  • 146865 06/22/05 01:37 Markowitz adaql3 hung, using disks on /adaql1/data2
  • 146863 06/22/05 00:47 Raue Septum IOC
  • 146857 06/22/05 00:09 breuer shift summary
  • Tuesday

  • 146851 06/21/05 22:12 breuer sli screen, acc. screen
  • 146850 06/21/05 22:09 breuer rich strip chart and hv
  • 146849 06/21/05 22:07 breuer right septum strip chart
  • 146848 06/21/05 22:07 breuer magnet level strip charts
  • 146847 06/21/05 22:06 breuer beam current strip charts
  • 146846 06/21/05 22:03 breuer general tools screen
  • 146843 06/21/05 21:34 aniol general tools
  • 146842 06/21/05 21:33 aniol rich screen
  • 146827 06/21/05 16:04 C. C. Chang Shift summary
  • 146824 06/21/05 15:44 C. C. Chang No beam
  • 146823 06/21/05 15:06 C. C. Chang No beam
  • 146820 06/21/05 14:37 C. C. Chang SLI screen shot
  • 146819 06/21/05 14:33 C. C. Chang RICH strip charts
  • 146818 06/21/05 14:32 C. C. Chang Right septum strip charts
  • 146817 06/21/05 14:29 C. C. Chang Magnet level strip charts
  • 146816 06/21/05 14:26 C. C. Chang Hall A General Tools
  • 146811 06/21/05 12:29 C. C. Chang Right septa screen shot
  • 146810 06/21/05 12:19 C. C. Chang Right Septum magnet trip
  • 146808 06/21/05 11:43 Iodice 16NLambda hyper spectrum
  • 146805 06/21/05 10:46 Markowitz correlation of bore shield and coil temperatures
  • 146798 06/21/05 08:46 C. C. Chang SLI/Beam energy spread
  • 146795 06/21/05 08:32 C. C. Chang SLI screen
  • 146794 06/21/05 08:19 C. C. Chang Right Q1 flow rate alarm
  • 146793 06/21/05 08:01 Raue Rich moisture
  • 146789 06/21/05 07:58 cussano/raue shift summary
  • 146788 06/21/05 07:57 A. Acha T/R RICH
  • 146787 06/21/05 07:55 Raue SLI screen
  • 146782 06/21/05 05:53 Raue septum trip
  • 146779 06/21/05 04:48 A. Acha New reference plots
  • 146776 06/21/05 04:03 cusanno SLI screen shot
  • 146775 06/21/05 04:01 cusanno SLI screen shot
  • 146774 06/21/05 04:00 cusanno BCM/SLI snapshot
  • 146771 06/21/05 03:14 cusanno Beam back at 03:00
  • 146768 06/21/05 02:47 cusanno BCM/SLI stripchart screenshot
  • 146767 06/21/05 02:34 cusanno Beam off to put Hall C viewer
  • 146766 06/21/05 02:10 cusanno SLI screenshot
  • 146765 06/21/05 02:08 cusanno BCM screenshot
  • 146764 06/21/05 01:57 Q1 cryo flow: techs
  • 146760 06/21/05 00:53 cusanno SLI real screen shot
  • 146758 06/21/05 00:44 cusanno SLI screen shot
  • 146757 06/21/05 00:32 breuer shift summary
  • 146755 06/21/05 00:14 Markowitz runs 4758--4765 junk (DAQ problem)
  • Monday

  • 146735 06/20/05 21:34 ellis sli screen
  • 146734 06/20/05 21:27 ellis right_septum.stp
  • 146733 06/20/05 21:26 ellis magneticlevels1.stp, magneticlevels2.stp
  • 146732 06/20/05 21:24 ellis dipole-septum.stp
  • 146731 06/20/05 21:22 breuer bcm_spread, dipole-septum strip chart
  • 146730 06/20/05 21:20 breuer general tools screen, beam current strip chart
  • 146718 06/20/05 17:30 A. Acha pmt adjustment A1
  • 146716 06/20/05 16:52 A. Acha PMT volatge change in A1
  • 146713 06/20/05 16:12 LeRose Right Septum Trips
  • 146711 06/20/05 16:08 LeRose BPM's back
  • 146710 06/20/05 16:07 A. Acha replaced pmt A1
  • 146709 06/20/05 16:06 LeRose Shift Summary
  • 146706 06/20/05 15:46 LeRose BPM's missing
  • 146703 06/20/05 15:15 LeRose BCM/SLI
  • 146697 06/20/05 13:58 LeRose SLI
  • 146696 06/20/05 13:33 LeRose A1 Channel 10 HV
  • 146692 06/20/05 12:21 LeRose Tools
  • 146691 06/20/05 12:18 LeRose SLI
  • 146689 06/20/05 11:49 LeRose Beam Back
  • 146688 06/20/05 11:13 LeRose Moller Target
  • 146687 06/20/05 11:12 Markowitz A1 pmt replaced
  • 146686 06/20/05 10:49 LeRose Moller Target is out
  • 146685 06/20/05 10:29 LeRose Right Q1 Lead Flow
  • 146684 06/20/05 10:25 Ewing Techs
  • 146683 06/20/05 08:34 LeRose Beam Off
  • 146680 06/20/05 07:59 Raue Shift summary
  • 146676 06/20/05 06:33 Raue Septum trip
  • 146674 06/20/05 06:10 Raue SLI screen
  • 146673 06/20/05 06:09 Raue good running
  • 146668 06/20/05 04:28 Raue printer
  • 146663 06/20/05 03:07 raue sli screen
  • 146651 06/20/05 00:04 breuer shift summary 16:00-24:00
  • Sunday

  • 146648 06/19/05 23:15 breuer copy of target window
  • 146647 06/19/05 23:14 breuer copy of Rich detector page
  • 146646 06/19/05 23:13 breuer right septum strip chart
  • 146641 06/19/05 21:42 breuer magnet strip charts
  • 146640 06/19/05 21:40 breuer BCM strip chart
  • 146639 06/19/05 21:39 breuer general tools page, BCM
  • 146631 06/19/05 19:30 breuer right septum trip, no trip
  • 146626 06/19/05 17:18 Markowitz Kaon coincidence timing
  • 146624 06/19/05 16:37 R. Michaels test of elog.php
  • 146623 06/19/05 16:29
  • 146620 06/19/05 15:56 E. Cisbani Shift Summary (Day 06/19)
  • 146618 06/19/05 15:51
  • 146615 06/19/05 15:46 E. Cisbani BCM screen
  • 146614 06/19/05 15:45 E. Cisbani Hall A Tools screen
  • 146613 06/19/05 15:16
  • 146612 06/19/05 15:09 E. Cisbani SLI screen
  • 146610 06/19/05 14:54
  • 146609 06/19/05 14:52 E. Cisbani Water Speed reduced to 600
  • 146607 06/19/05 14:19
  • 146606 06/19/05 14:04
  • 146603 06/19/05 13:49 E. Cisbani SLI screen
  • 146602 06/19/05 13:48 E. Cisbani Right Septum Upper Coil Over Temperature
  • 146599 06/19/05 13:27
  • 146595 06/19/05 13:06
  • 146594 06/19/05 12:36 Cisbani Right Septum trip
  • 146593 06/19/05 12:31
  • 146591 06/19/05 12:15
  • 146590 06/19/05 11:55 E. Cisbani SLI
  • 146587 06/19/05 11:34
  • 146586 06/19/05 11:20
  • 146583 06/19/05 10:45 cusanno RICH event display
  • 146582 06/19/05 10:44
  • 146579 06/19/05 10:33 E.Cisbani SLI screen
  • 146578 06/19/05 10:22
  • 146575 06/19/05 09:42
  • 146574 06/19/05 09:40
  • 146571 06/19/05 09:28 E. Cisbani SLI screen
  • 146570 06/19/05 09:15 Markowitz adaql2 reboot / CODA trouble
  • 146568 06/19/05 08:47
  • 146567 06/19/05 08:47
  • 146566 06/19/05 08:03 V. Sulkosky SLI and BCM Stripchart
  • 146564 06/19/05 08:00 Raue Shift summary
  • 146563 06/19/05 07:52
  • 146562 06/19/05 07:52
  • 146561 06/19/05 07:04 V. Sulkosky SLI screen
  • 146560 06/19/05 06:59
  • 146559 06/19/05 06:58
  • 146558 06/19/05 06:15 V. Sulkosky Problem with Automatic Halog Entries
  • 146557 06/19/05 06:04
  • 146556 06/19/05 06:04
  • 146555 06/19/05 06:01 V. Sulkosky Hall A Tools
  • 146554 06/19/05 06:00 V. Sulkosky RICH screens
  • 146553 06/19/05 05:54 V. Sulkosky SLI screen
  • 146552 06/19/05 05:12
  • 146551 06/19/05 05:12
  • 146550 06/19/05 04:15
  • 146549 06/19/05 04:15 V. Sulkosky SLI screen
  • 146548 06/19/05 04:10
  • 146545 06/19/05 03:33
  • 146544 06/19/05 03:14
  • 146541 06/19/05 03:02 raue sl1 snapshot
  • 146540 06/19/05 02:47 L. Lagamba Rich detector (Number of hits plot)
  • 146539 06/19/05 02:37
  • 146537 06/19/05 02:22
  • 146535 06/19/05 01:46
  • 146533 06/19/05 01:28
  • 146532 06/19/05 01:07 Markowitz ion chamber trips and vacuum valve closings
  • 146531 06/19/05 01:00 Markowitz A1 aerogel PMT
  • 146530 06/19/05 00:52
  • 146529 06/19/05 00:29 L. Lagamba Rich detector (Number of cluster)
  • 146528 06/19/05 00:05 breuer RICH detector screen
  • Saturday

  • 146526 06/18/05 23:56 Yi Qiang Target Snapshot
  • 146525 06/18/05 23:55 breuer shift summary 16:00 - 24:00
  • 146524 06/18/05 23:41
  • 146523 06/18/05 23:26
  • 146521 06/18/05 22:05
  • 146520 06/18/05 22:04
  • 146513 06/18/05 20:42 breuer SLI SCREEN (2.8%)
  • 146512 06/18/05 20:37 breuer RICH strip chart and HV
  • 146511 06/18/05 20:37
  • 146510 06/18/05 20:34 breuer General Tools screen, septum strip ch.
  • 146509 06/18/05 20:31 breuer strip charts: BCM
  • 146508 06/18/05 20:22
  • 146502 06/18/05 19:29 F. Garibaldi Shift Summary
  • 146501 06/18/05 19:08
  • 146498 06/18/05 18:49
  • 146495 06/18/05 17:47 markowitz mcc lost arc 1 box supply
  • 146493 06/18/05 17:37
  • 146490 06/18/05 17:11 markowitz right q1, helium level: techs
  • 146489 06/18/05 17:02
  • 146488 06/18/05 17:00 markowitz rich
  • 146486 06/18/05 16:35 Yi Qiang Right Septum Tripped
  • 146485 06/18/05 16:27 cusanno Lambda production on last two days
  • 146481 06/18/05 16:07
  • 146477 06/18/05 15:19
  • 146476 06/18/05 15:05 Garibaldi septa strip chart
  • 146473 06/18/05 14:43
  • 146472 06/18/05 14:37 F. Garibaldi SLI
  • 146471 06/18/05 14:19 E. Cisbani SLI
  • 146466 06/18/05 13:34
  • 146463 06/18/05 13:13
  • 146462 06/18/05 13:04 F. Garibaldi SLI
  • 146458 06/18/05 12:01
  • 146455 06/18/05 11:45 F. Garibaldi SLI
  • 146454 06/18/05 11:43
  • 146446 06/18/05 10:49 F. Garibaldi SLI
  • 146445 06/18/05 10:16
  • 146442 06/18/05 10:12
  • 146437 06/18/05 09:00 Markowitz techs: right septum temperature
  • 146436 06/18/05 08:30 F. Garibaldi Lowering beam current
  • 146433 06/18/05 08:17
  • 146432 06/18/05 08:17
  • 146430 06/18/05 08:05 V. Sulkosky Owl Shift Summary
  • 146427 06/18/05 07:51 V. Sulkosky SLI screen
  • 146425 06/18/05 07:24 V. Sulkosky SLI and BCM charts
  • 146421 06/18/05 06:47
  • 146420 06/18/05 06:41 V. Sulkosky SLI screen
  • 146419 06/18/05 06:30
  • 146418 06/18/05 06:23 V. Sulkosky Hall A Tools
  • 146416 06/18/05 05:47 V. Sulkosky Septum stripchart
  • 146415 06/18/05 05:33 V. Sulkosky beam gone
  • 146413 06/18/05 05:16
  • 146412 06/18/05 04:57
  • 146409 06/18/05 04:40 V. Sulkosky SLI screen
  • 146406 06/18/05 03:49 V. Sulkosky Right septum temp fluctuations
  • 146403 06/18/05 03:24 V. Sulkosky Septum Strip chart
  • 146402 06/18/05 03:09 V. Sulkosky SLI screen
  • 146398 06/18/05 02:54 L. Lagamba Physics Replay code
  • 146397 06/18/05 02:43
  • 146396 06/18/05 02:35 V. Sulkosky Right septum up after second trip!
  • 146394 06/18/05 01:56 V. Sulkosky Right septum trip
  • 146393 06/18/05 01:52
  • 146391 06/18/05 01:45 V. Sulkosky SLI screen
  • 146390 06/18/05 01:22 L. Lagamba RICH detector
  • 146387 06/18/05 00:45 l. Lagamba Rich detector
  • 146384 06/18/05 00:28 V. Sulkosky SLI screen
  • 146383 06/18/05 00:07 E.Chudakov Swing shift summary
  • 146381 06/18/05 00:03
  • 146380 06/18/05 00:01 E.Chudakov Strip charts
  • Friday

  • 146376 06/17/05 22:49
  • 146373 06/17/05 22:37 E.Chudakov SLI screen
  • 146368 06/17/05 21:36
  • 146367 06/17/05 21:24 E.Chudakov SLI screen
  • 146362 06/17/05 20:25 E.Chudakov SLI screen
  • 146361 06/17/05 20:23 E,Chudakov Strip charts
  • 146358 06/17/05 19:47
  • 146354 06/17/05 19:07
  • 146353 06/17/05 18:58 E.Chudakov RICH HV tripped
  • 146346 06/17/05 17:26 E.Chudakov Septa charts
  • 146345 06/17/05 17:23 E.Chudakov Septa are back
  • 146343 06/17/05 17:19 E.Chudakov SLI screen
  • 146342 06/17/05 17:10 E.Chudakov Both septa tripped at 16:05
  • 146341 06/17/05 16:43
  • 146340 06/17/05 16:29
  • 146337 06/17/05 16:01 LeRose Shift Summary
  • 146334 06/17/05 15:40 LeRose Hall A Tools
  • 146333 06/17/05 15:31 A. Acha runs
  • 146330 06/17/05 14:58 LeRose BCM Chart
  • 146327 06/17/05 14:34 LeRose SLI Snapshot
  • 146325 06/17/05 14:01
  • 146324 06/17/05 13:58 A. Acha Good Runs
  • 146323 06/17/05 13:51 LeRose Right Septum Trip
  • 146322 06/17/05 13:26
  • 146320 06/17/05 13:02 LeRose Vacuum Valve Trip
  • 146319 06/17/05 12:52 LeRose SLI Snapshot
  • 146314 06/17/05 11:37 LeRose Good Beam
  • 146313 06/17/05 11:20
  • 146310 06/17/05 10:49 LeRose SLI Snapshot
  • 146307 06/17/05 10:25 A. Acha Rich trips
  • 146305 06/17/05 10:13 LeRose RICH Trips again!
  • 146304 06/17/05 10:10
  • 146301 06/17/05 10:05 LeRose RICH Trip
  • 146298 06/17/05 09:45 LeRose RICH Trip
  • 146294 06/17/05 09:03 LeRose SLI Snapshot
  • 146293 06/17/05 08:45 LeRose RICH Trip
  • 146290 06/17/05 08:38
  • 146288 06/17/05 07:56 R. Michaels owl summary
  • 146286 06/17/05 07:54 R. Michaels RICH HV trip - end run 4620
  • 146284 06/17/05 07:49 cusanno RICH is working fine
  • 146283 06/17/05 07:32 R. Michaels SLI ok again ?
  • 146281 06/17/05 07:20 R. Michaels 8hr BCM & slia
  • 146279 06/17/05 07:07
  • 146276 06/17/05 06:35 R. Michaels beam centering
  • 146274 06/17/05 06:15 R. Michaels SLI "not working", run anyway
  • 146273 06/17/05 06:03 R. Michaels screen (RICH & Tools)
  • 146271 06/17/05 05:39 R. Michaels beam back
  • 146269 06/17/05 04:08 luigi detector check
  • 146268 06/17/05 03:52
  • 146267 06/17/05 03:37 R. Michaels beam down -- PSS problems
  • 146264 06/17/05 03:22 R. Michaels beam returns, work on E res.
  • 146263 06/17/05 03:06 Luigi Detector checks (RICH cluster?)
  • 146261 06/17/05 01:14
  • 146258 06/17/05 00:59 R. Michaels status update
  • 146252 06/17/05 00:09 jpchen Swing shift summary
  • 146251 06/17/05 00:06 jpchen RICH status and another RICH PED run
  • Thursday

  • 146248 06/16/05 23:50 jpchen hall a screen and BCM
  • 146247 06/16/05 23:44 jpchen RICH control settings
  • 146246 06/16/05 23:41 jpchen RICH spectra seem not right
  • 146245 06/16/05 23:30 Markowitz A1 channel 10
  • 146244 06/16/05 23:27 shzhou/jpche A1 ADC10 low
  • 146243 06/16/05 23:06
  • 146239 06/16/05 21:44 jpchen Pulsed beam back
  • 146238 06/16/05 21:10 R. Feuerbach To reboot hallavme17 (ROC7)
  • 146237 06/16/05 19:57 jpchen RICH HV ch16 back to 2120V
  • 146236 06/16/05 19:50 folts tech
  • 146235 06/16/05 19:50 jpchen Q2/Q3 cycled, magnets ready
  • 146234 06/16/05 19:36 Markowitz All magnets OK
  • 146233 06/16/05 19:24
  • 146230 06/16/05 18:10
  • 146229 06/16/05 17:18 apar End_of_Parity_Run_3448__CODA_Config=happex2_min
  • 146228 06/16/05 17:14 apar Start_Parity_Run_3448__CODA_Config=happex2_min
  • 146227 06/16/05 17:14 apar End_of_Parity_Run_3447__CODA_Config=happex2_min
  • 146226 06/16/05 17:11 apar Start_Parity_Run_3447__CODA_Config=happex2_min
  • 146225 06/16/05 16:14 C. C. Chang Shift summary
  • 146224 06/16/05 16:13 R. Feuerbach Plan for this evening
  • 146223 06/16/05 16:10 F. Cusanno, printouts of analyses with RICH
  • 146221 06/16/05 13:37
  • 146220 06/16/05 13:19
  • 146219 06/16/05 08:12 C. C. Chang RICH HV
  • 146218 06/16/05 07:56 M.Iodice Owl shift summary
  • 146217 06/16/05 07:14
  • 146216 06/16/05 03:45
  • 146215 06/16/05 02:42 Iodice Waterfall density stability tests
  • 146214 06/16/05 02:41 apar End_of_Parity_Run_3446__CODA_Config=happex2_min
  • 146211 06/16/05 02:35 apar Start_Parity_Run_3446__CODA_Config=happex2_min
  • 146209 06/16/05 02:34 apar End_of_Parity_Run_3445__CODA_Config=happex2_min
  • 146207 06/16/05 02:27 apar Start_Parity_Run_3445__CODA_Config=happex2_min
  • 146205 06/16/05 02:26 apar End_of_Parity_Run_3444__CODA_Config=happex2_min
  • 146203 06/16/05 02:22 apar Start_Parity_Run_3444__CODA_Config=happex2_min
  • 146201 06/16/05 02:20 apar End_of_Parity_Run_3443__CODA_Config=happex2_min
  • 146199 06/16/05 02:14 apar Start_Parity_Run_3443__CODA_Config=happex2_min
  • 146196 06/16/05 01:59 apar End_of_Parity_Run_3442__CODA_Config=happex2_min
  • 146194 06/16/05 01:54 apar Start_Parity_Run_3442__CODA_Config=happex2_min
  • 146193 06/16/05 01:03 E.Chudakov Swing shift summary
  • 146191 06/16/05 00:30 cusanno Preliminary Oxygen missing energy spectrum
  • Wednesday

  • 146190 06/15/05 23:47 V. Sulkosky RICH HV ON, flow to 800
  • 146189 06/15/05 23:43 V. Sulkosky Blumi HVs Set
  • 146188 06/15/05 22:29
  • 146186 06/15/05 18:38 folts tech
  • 146185 06/15/05 18:30 folts tech
  • 146184 06/15/05 17:53 V. Sulkosky Left and Right Q1's off
  • 146183 06/15/05 17:47 V. Sulkosky Left Q1 ramping down
  • 146182 06/15/05 17:22
  • 146181 06/15/05 16:10 R. Feuerbach Plan for this evening
  • 146180 06/15/05 16:09 C. C. Chang Right Q1 trip
  • 146178 06/15/05 16:04 Higinbotham More Plots (zoomed)
  • 146177 06/15/05 15:48 Higinbotham Ta Plots from Run 4277
  • 146176 06/15/05 15:20 C. C. Chang Right Q3
  • 146175 06/15/05 14:52 C. C. Chang Right Q3 ramping
  • 146174 06/15/05 13:53
  • 146173 06/15/05 12:39 cusanno Coincidence Time spectrum, RICH in
  • 146172 06/15/05 11:46 Jack Left Arm Q2
  • 146171 06/15/05 11:15 C. C. Chang Right Q3 ramping down
  • 146170 06/15/05 10:20 Reitz bringing up the RICH
  • 146169 06/15/05 08:56 X. Zheng Cherenkov ana sum - PV DIS test
  • 146168 06/15/05 08:36 C. C. Chang Left Q2 trip
  • 146167 06/15/05 05:29
  • 146166 06/15/05 02:56 E.Chudakov No Moller test
  • Tuesday

  • 146164 06/14/05 22:09 R. Feuerbach Shifts cancelled/standby
  • 146163 06/14/05 22:07
  • 146162 06/14/05 21:56 V. Sulkosky Turned Lumi HV off
  • 146161 06/14/05 21:49 Eric Voutier Shift Summary
  • 146160 06/14/05 21:43 V. Sulkosky Hall A Tools during Density Tests
  • 146159 06/14/05 21:41 V. Sulkosky Turn RICH HV Off and Flow to 410
  • 146158 06/14/05 21:37 V. Sulkosky Move Target to Empty
  • 146156 06/14/05 21:31 apar End_of_Parity_Run_3441__CODA_Config=happex2_min
  • 146154 06/14/05 21:27 apar Start_Parity_Run_3441__CODA_Config=happex2_min
  • 146153 06/14/05 21:25 apar End_of_Parity_Run_3440__CODA_Config=happex2_min
  • 146150 06/14/05 21:15 apar Start_Parity_Run_3440__CODA_Config=happex2_min
  • 146148 06/14/05 21:13 apar End_of_Parity_Run_3439__CODA_Config=happex2_min
  • 146146 06/14/05 21:06 apar Start_Parity_Run_3439__CODA_Config=happex2_min
  • 146144 06/14/05 21:05 apar End_of_Parity_Run_3438__CODA_Config=happex2_min
  • 146142 06/14/05 20:59 apar Start_Parity_Run_3438__CODA_Config=happex2_min
  • 146140 06/14/05 20:57 apar End_of_Parity_Run_3437__CODA_Config=happex2_min
  • 146138 06/14/05 20:51 apar Start_Parity_Run_3437__CODA_Config=happex2_min
  • 146137 06/14/05 20:39 V. Sulkosky Blumi HV set
  • 146136 06/14/05 20:09 R. Feuerbach Plan for tonight
  • 146135 06/14/05 19:53 Fansler Techs
  • 146134 06/14/05 18:34 V. Sulkosky Left Q3 Problems
  • 146133 06/14/05 16:46 Fansler Techs
  • 146132 06/14/05 16:10 C. C. chang Shift summary
  • 146130 06/14/05 15:53 C. C. Chang SLI
  • 146129 06/14/05 15:43 C. C. Chang RICH strip charts
  • 146128 06/14/05 15:41 C. C. Chang Septa screenshots
  • 146127 06/14/05 15:39 C. C. Chang Septa strip charts
  • 146126 06/14/05 15:36 C. C. Chang Septa dipole strip charts
  • 146125 06/14/05 15:35 C. C. Chang Magnet strip charts
  • 146124 06/14/05 15:25 C. C. Chang Hall A screenshot
  • 146123 06/14/05 13:14
  • 146122 06/14/05 12:08 apar End_of_Parity_Run_3436__CODA_Config=happex2_min
  • 146121 06/14/05 11:53 apar Start_Parity_Run_3436__CODA_Config=happex2_min
  • 146120 06/14/05 11:20 C. C. Chang Magnet strip charts
  • 146119 06/14/05 11:14 C. C. Chang Magnets trip
  • 146118 06/14/05 10:52 R. Feuerbach RICH gas swapped
  • 146117 06/14/05 10:16 R. Feuerbach Real problems with R-Q1 power supply
  • 146116 06/14/05 09:58 C. C. Chang Septa strip charts
  • 146115 06/14/05 09:51 C. C. Chang Right Q1 problem
  • 146114 06/14/05 09:49 R. Feuerbach INSTRUCTIONS: Monitoring SLI
  • 146109 06/14/05 07:58 E.Chudakov Owl shift summary
  • 146106 06/14/05 07:08 E.Chudakov RICH HV - back to 2120V
  • 146101 06/14/05 06:23
  • 146100 06/14/05 06:07 E.Chudakov Beam current strip chart
  • 146099 06/14/05 06:04 E.Chudakov RICH HV = 2080V
  • 146094 06/14/05 04:55 E.Chudakov RICH HV = 2100V
  • 146089 06/14/05 02:48 E.Chudakov The arc magnet tripped again
  • 146088 06/14/05 02:38 E.Chudakov Disk problems
  • 146087 06/14/05 02:25 E.Chudakov Data analysis - disk problem
  • 146084 06/14/05 02:14 E.Chudakov Beam was off for 30 min due to a magnet in the machine
  • 146083 06/14/05 01:12 E.Chudakov Energy spread
  • 146082 06/14/05 01:07 E.Chudakov Going to 80uA
  • 146079 06/14/05 00:03 Reitz RICH Commissioning
  • 146078 06/14/05 00:03 Perdrisat Shift record
  • Monday

  • 146072 06/13/05 23:36 R. Feuerbach OTR and SLI
  • 146070 06/13/05 23:21 cusanno RICH ring
  • 146062 06/13/05 22:35 V. Sulkosky Parity Runs II
  • 146053 06/13/05 21:37 V. Sulkosky SLI screen
  • 146052 06/13/05 21:23 V. Sulkosky Septum Controls
  • 146051 06/13/05 21:21 V. Sulkosky Current and Energy spread scriptcharts
  • 146050 06/13/05 21:13 V. Sulkosky Hall A General Tools
  • 146048 06/13/05 20:43 Markowitz Updated database beam energy, after EP measurement
  • 146047 06/13/05 19:58 Markowitz Waterfall coincidence timing and missing mass
  • 146046 06/13/05 19:47 apar End_of_Parity_Run_3435__CODA_Config=happex2_min
  • 146044 06/13/05 19:35 Reitz EP measurement
  • 146043 06/13/05 19:35 apar Start_Parity_Run_3435__CODA_Config=happex2_min
  • 146041 06/13/05 18:43 V. Sulkosky Run: 4571 stopped for ep Measurement
  • 146040 06/13/05 18:41 V. Sulkosky Parity runs I
  • 146039 06/13/05 18:27 apar End_of_Parity_Run_3434__CODA_Config=happex2_min
  • 146037 06/13/05 18:21 apar Start_Parity_Run_3434__CODA_Config=happex2_min
  • 146035 06/13/05 18:16 apar End_of_Parity_Run_3433__CODA_Config=happex2_min
  • 146033 06/13/05 17:42 apar Start_Parity_Run_3433__CODA_Config=happex2_min
  • 146025 06/13/05 16:20 LeRose Left Arm Controls reboot
  • 146023 06/13/05 16:04 LeRose Shift Summary
  • 146022 06/13/05 16:00 LeRose BPM Controlled Access
  • 146020 06/13/05 15:48 A. Acha, R. Comparison of 500 and 750 Tach signal/background
  • 146015 06/13/05 13:49 LeRose Back in Business
  • 146014 06/13/05 13:46 LeRose Right Septum Returning
  • 146013 06/13/05 13:28 LeRose Right Septum Too
  • 146011 06/13/05 13:18 LeRose Cryogenic Event TECH
  • 146006 06/13/05 12:13 LeRose RICH installed
  • 146005 06/13/05 11:50 Jack RICH installation
  • 146004 06/13/05 11:01
  • 146003 06/13/05 10:27 LeRose RICH is Back
  • 146002 06/13/05 09:06 Higinbotham To Keep Sieve Problem From Happening Again
  • 146000 06/13/05 08:13 cusanno waterfall CT spectrum is file at 50 uA, Tach 750 too
  • 145999 06/13/05 08:07 Guido Shift Summary
  • 145994 06/13/05 04:49
  • 145992 06/13/05 04:15 G.M Urciuoli OTR and SLI
  • 145982 06/13/05 00:59 cusanno CT time is good with Waterfall
  • 145979 06/13/05 00:08 cusanno Swing Shift Summary
  • 145978 06/13/05 00:01 cusanno Tach to 500 for the next two runs
  • 145977 06/13/05 00:00 cusanno prescaler change
  • Sunday

  • 145976 06/12/05 23:59 cusanno Beam off
  • 145974 06/12/05 23:49 cusanno Septum strip chart
  • 145973 06/12/05 23:48 cusanno Hall A screenshot
  • 145972 06/12/05 23:22 Ramesh Y_target puzzle of right hrs solved
  • 145970 06/12/05 22:59 R. Feuerbach R-HRS Sieve partially in
  • 145969 06/12/05 22:43 Ramesh Controlled access requested
  • 145967 06/12/05 22:23 R. Feuerbach Tools snapshot for test
  • 145966 06/12/05 22:16 R. Subedi, F Troublshooting of right hrs y_target value
  • 145958 06/12/05 19:31 Adaq halla tools
  • 145954 06/12/05 18:07 cusanno Snapshot of OTR
  • 145952 06/12/05 17:48 R. Michaels re: efficiencies
  • 145951 06/12/05 17:22 R. Michaels Run 4502 cleanup
  • 145950 06/12/05 17:17 R. Feuerbach Re: "problems" with VDCs
  • 145949 06/12/05 17:10 Urciuoli Day Shift Summary
  • 145945 06/12/05 16:16 cusanno SLI screenshot
  • 145944 06/12/05 15:24 S. Marrone Run 4542 is with C-target
  • 145943 06/12/05 15:22 S. Marrone VDC problem
  • 145941 06/12/05 15:20 S. Marrone Change to Waterfall target
  • 145939 06/12/05 13:26 R. Michaels coinc. timing is ok
  • 145938 06/12/05 12:56 Markowitz Signal-to-Noise ratio, carbon and waterfall
  • 145937 06/12/05 12:07 G.M. Urciuol Target Change
  • 145935 06/12/05 11:30 S. Marrone VDC L-arm problem
  • 145932 06/12/05 10:22 R. Feuerbach OTR and SLI screens.
  • 145931 06/12/05 10:08 G.M. Urciuol SLI
  • 145930 06/12/05 10:06 G.M. Urciuol OTR
  • 145929 06/12/05 08:06 Boeglin Shift Summary
  • 145927 06/12/05 07:55 Adaq Strip chart snap shot
  • 145926 06/12/05 07:51 Boeglin Magnets ramping up again
  • 145925 06/12/05 07:24 Boeglin Magnet Trips
  • 145921 06/12/05 06:13 Boeglin SLI screen shot
  • 145920 06/12/05 06:12 Boeglin OTR screen shot
  • 145911 06/12/05 03:57 Boeglin SLI screen shot
  • 145906 06/12/05 02:50 Boeglin SLI screann shot after beam returns, run 4533
  • Saturday

  • 145903 06/11/05 23:54 S. Marrone Shift Summary
  • 145901 06/11/05 21:52 adaq SLI
  • 145894 06/11/05 20:26 cusanno PS5 set to 20
  • 145893 06/11/05 20:17 adaq BCM chart
  • 145890 06/11/05 19:59 adaq hall a magnets
  • 145889 06/11/05 19:49 adaq SLI
  • 145882 06/11/05 18:15 R. Michaels T7 efficiency with waterfall
  • 145880 06/11/05 17:11 Adaq SLI
  • 145878 06/11/05 16:30
  • 145876 06/11/05 16:26 G.M. Urciuol Day shift report
  • 145875 06/11/05 16:15 Adaq SLI
  • 145874 06/11/05 16:09 adaq OTR stuff
  • 145860 06/11/05 14:00 Markowitz tacheometer to 750
  • 145857 06/11/05 13:45 Fansler Techs
  • 145856 06/11/05 13:24 Jack RICH is out of Left Arm
  • 145855 06/11/05 11:40 Lindgren Tec - Q1 Right trip
  • 145854 06/11/05 11:03
  • 145853 06/11/05 09:54 Lindgren Hall A access to remove RICH
  • 145852 06/11/05 09:49 Reitz RICH Freon Leak
  • 145844 06/11/05 08:51 Adaq right Q1 tripped; techs
  • 145839 06/11/05 08:04 Boeglin Shift Summary
  • 145836 06/11/05 07:22 Adaq BCM etc. strip chart
  • 145835 06/11/05 07:19 Boeglin OTR screen shot
  • 145834 06/11/05 06:52 boeglin SLI screen shot
  • 145833 06/11/05 06:45 Boeglin All magnets are up again
  • 145829 06/11/05 04:52 Boeglin Right arm dipole ramping up
  • 145828 06/11/05 04:50 Boeglin Right Septum recovered
  • 145827 06/11/05 04:45 Boeglin Right Septum recovered
  • 145826 06/11/05 03:51 cusanno Waterfall spectrum
  • 145825 06/11/05 03:47 cusanno Waterfall spectrum
  • 145824 06/11/05 03:30 Boeglin Spectrometer trips, Septum trips
  • 145822 06/11/05 03:00 boeglin SLI snap shot
  • 145820 06/11/05 01:40 Boeglin RICH instructions for the night
  • 145819 06/11/05 01:26 C. Perdrisat Shift report
  • 145817 06/11/05 01:05 Boeglin Problem with run 4502-4503
  • Friday

  • 145802 06/10/05 23:49 E.Chudakov Shower/Preshower calibrated
  • 145801 06/10/05 23:48 a-onl /adaqfs/home/a-onl/epics/runfiles/halog_end_4502.epics
  • 145799 06/10/05 23:37 a-onl /adaqfs/home/a-onl/epics/runfiles/halog_end_4502.epics
  • 145798 06/10/05 23:33 a-onl /adaqfs/home/a-onl/epics/runfiles/halog_end_4502.epics
  • 145768 06/10/05 16:53 spiegel Hall A tech's
  • 145765 06/10/05 16:06 G.M. Urciuol Day shift report
  • 145760 06/10/05 13:32 reitz hall a access
  • 145759 06/10/05 13:30 Reitz parasitic rich commissioning
  • 145744 06/10/05 09:54 apar End_of_Parity_Run_3432__CODA_Config=Lumi
  • 145743 06/10/05 09:52 apar Start_Parity_Run_3432__CODA_Config=Lumi
  • 145742 06/10/05 09:51 apar End_of_Parity_Run_3431__CODA_Config=Lumi
  • 145741 06/10/05 09:50 apar Start_Parity_Run_3431__CODA_Config=Lumi
  • 145740 06/10/05 09:32 R. Feuerbach SLI is INDEPENDENT of Beam current
  • 145738 06/10/05 08:59 R. Feuerbach SLI vs OTR beam sizes
  • 145733 06/10/05 08:05 Boeglin Shift Summary (recreated)
  • 145731 06/10/05 07:53 R. Michaels BPM calibration problem ?
  • 145730 06/10/05 07:51 boeglin SLI strip chart snap shot
  • 145729 06/10/05 07:49 boeglin SLI screen shot
  • 145724 06/10/05 06:24 boeglin SLI strip chart runs 4472, 4473
  • 145723 06/10/05 06:18 R. Michaels adaql1 reboot (again!)
  • 145719 06/10/05 05:42 R. Michaels deadtime vs trigger
  • 145718 06/10/05 05:36 R. Michaels Start up again
  • 145714 06/10/05 05:28 boeglin sli 25 uA
  • 145713 06/10/05 05:27 boeglin otr 25 uA
  • 145712 06/10/05 05:26 boeglin otr 100 uA
  • 145711 06/10/05 05:25 boeglin sli 100 uA
  • 145710 06/10/05 05:24 boeglin sli 75 uA with otr
  • 145709 06/10/05 05:23 boeglin otr 75uA
  • 145708 06/10/05 04:59 Boeglin otr 50 uA
  • 145707 06/10/05 04:54 boeglin sli 50uA with otr
  • 145706 06/10/05 04:53 boeglin sli 50uA
  • 145705 06/10/05 04:01 R. Michaels Re-enabled event type 100
  • 145702 06/10/05 02:56 Boeglin left arm magnet control after reboot of iocha14
  • 145701 06/10/05 02:37 Boeglin reboot iocha14, left arm magnet control
  • 145700 06/10/05 02:01 R. Michaels PS file for cosmics / status
  • 145697 06/10/05 00:20 R. Michaels deadtimes ok
  • 145695 06/10/05 00:06 Eric Voutier Shift Summary
  • Thursday

  • 145692 06/09/05 23:39 cusanno T7 efficiency and Kaon Identitication (no RICH)
  • 145674 06/09/05 23:14 Eric Voutier Tech / Q1-Left pb
  • 145647 06/09/05 21:47 E.Chudakov Right arm shower/preshower
  • 145645 06/09/05 21:39 reitz high deadtimes
  • 145637 06/09/05 20:05 Eric Voutier Tech / Q2 Left problem
  • 145629 06/09/05 17:12 R. Feuerbach Plan for 9-June through 10-June
  • 145628 06/09/05 16:04 Adaq Day shift report
  • 145626 06/09/05 10:47 LeRose Septum Setting is Now Automatic
  • 145625 06/09/05 08:39 Lindgren Hall A access
  • 145620 06/09/05 08:04 Boeglin Shift Summary
  • 145619 06/09/05 08:02 Boeglin BCM and Septum strip charts
  • 145618 06/09/05 08:01 Boeglin Main Screen Shot
  • 145615 06/09/05 05:32 cusanno Physics Replay
  • 145614 06/09/05 04:45 boeglin SLI snapshot
  • 145613 06/09/05 04:44 Adaq Strip chart snapshot
  • 145611 06/09/05 03:51 R. Feuerbach Updated S1 and S2 databases
  • 145603 06/09/05 00:30 R. Feuerbach Plan for the next day
  • Wednesday

  • 145601 06/08/05 23:52 sawatzky Swing shift summary
  • 145600 06/08/05 23:49 sawatzky BCM strip chart for shift
  • 145599 06/08/05 23:44 sawatzky General Tools, strip charts
  • 145598 06/08/05 23:40 sawatzky We must monitor the FFB status
  • 145595 06/08/05 23:15 E.Chudakov Left arm pion rejector is off
  • 145594 06/08/05 22:54 sawatzky SLI screen snap @ 100 uA
  • 145593 06/08/05 22:36 sawatzky General tools, strip charts
  • 145592 06/08/05 22:33 sawatzky Beam back, we asked to go to 100 uA (or best offer)
  • 145590 06/08/05 22:22 sawatzky Beam taken away for Hall C beam studies
  • 145588 06/08/05 22:04 sawatzky SLI screen snap
  • 145587 06/08/05 21:57 sawatzky Energy FFB turned on again
  • 145586 06/08/05 21:41 sawatzky General Tools, strip charts, beam looks good
  • 145585 06/08/05 21:37 sawatzky Prescales adjusted for 50 uA on Thick C
  • 145582 06/08/05 21:18 sawatzky Run 4425 is 50 uA production (w/ poor beam quality)
  • 145581 06/08/05 21:05 sawatzky Poor beam quality after the break
  • 145579 06/08/05 20:37 R. Michaels T7 efficiency (prelim.)
  • 145576 06/08/05 20:24 sawatzky Another MCC update
  • 145575 06/08/05 19:48 sawatzky MCC update
  • 145574 06/08/05 19:40
  • 145573 06/08/05 19:28 R. Michaels trigger timing check -- ok
  • 145572 06/08/05 17:40 sawatzky MCC update
  • 145571 06/08/05 17:17 sawatzky General Tools, Strip charts
  • 145570 06/08/05 16:58 Higinbotham Waterfall Y-Target
  • 145566 06/08/05 16:14 Higinbotham Updated Y-Target Matrix
  • 145565 06/08/05 16:06 Lindgren Shift summary
  • 145562 06/08/05 15:15 R. Michaels reboot CODA, start run 4421
  • 145560 06/08/05 15:10
  • 145555 06/08/05 14:51 R. Michaels Trigger time adjustment
  • 145554 06/08/05 14:50 atrig Trigger-Download
  • 0 06/08/05 14:09 no_name no_key
  • 0 06/08/05 13:43 no_name no_key
  • 145550 06/08/05 12:15 Lindgren Mid-Shift Summary
  • 145545 06/08/05 11:35 Lindgren Target change
  • 145543 06/08/05 11:31 apar End_of_Parity_Run_3430__CODA_Config=Lumi
  • 145542 06/08/05 11:29 apar Start_Parity_Run_3430__CODA_Config=Lumi
  • 145541 06/08/05 11:28 apar End_of_Parity_Run_3429__CODA_Config=Lumi
  • 145539 06/08/05 11:23 Lindgren Target Change
  • 145538 06/08/05 11:23 apar Start_Parity_Run_3429__CODA_Config=Lumi
  • 145534 06/08/05 10:55 apar End_of_Parity_Run_3428__CODA_Config=Lumi
  • 145533 06/08/05 10:53 apar Start_Parity_Run_3428__CODA_Config=Lumi
  • 145532 06/08/05 10:52 R. Feuerbach Scintillator trigger efficiency
  • 145531 06/08/05 10:51 Higinbotham BCM Problem Fixed
  • 145530 06/08/05 10:45 apar End_of_Parity_Run_3427__CODA_Config=Lumi
  • 145529 06/08/05 10:42 apar Start_Parity_Run_3427__CODA_Config=Lumi
  • 145526 06/08/05 10:31 apar End_of_Parity_Run_3426__CODA_Config=Lumi
  • 145525 06/08/05 10:30 apar Start_Parity_Run_3426__CODA_Config=Lumi
  • 145524 06/08/05 10:21 Lindgren Data acquisition configuration change
  • 145523 06/08/05 10:20 apar End_of_Parity_Run_3425__CODA_Config=Lumi
  • 145522 06/08/05 10:18 apar Start_Parity_Run_3425__CODA_Config=Lumi
  • 145521 06/08/05 10:18 Lindgren Beam back
  • 145520 06/08/05 10:16 Lindgren Beam off
  • 145515 06/08/05 10:03 Lindgren Beam back
  • 145514 06/08/05 09:59 Higinbotham Raster Image
  • 145511 06/08/05 09:29 Higinbotham Hall A General Tools
  • 145510 06/08/05 09:28 Higinbotham Beam Back SLI Image
  • 145509 06/08/05 09:17 Lindgren No beam.
  • 145508 06/08/05 09:05 R. Michaels ctime marginal
  • 145507 06/08/05 08:11 Adaq Shift Summary
  • 145504 06/08/05 07:21 Boeglin Changes to cuts_detReplay.def
  • 145503 06/08/05 07:21 Boeglin Changes to cuts_detReplay.def
  • 145499 06/08/05 06:29 Adaq Strip chart snapshot
  • 145495 06/08/05 05:08 Boeglin Problem with 100 uA running
  • 145492 06/08/05 04:22 Boeglin Strip chart snapshot
  • 145491 06/08/05 04:21 Boeglin SLI Snapshot at 50uA
  • 145484 06/08/05 02:39 Boeglin SLI snapshot @ 100 uA
  • 145481 06/08/05 02:33 cusanno Lamda peak
  • 145476 06/08/05 01:37 R. Feuerbach caution about CTOF distribution
  • 145474 06/08/05 01:23 Boeglin SLI strip chart snapshot
  • 145472 06/08/05 01:21 Boeglin SLI snapshot
  • 145467 06/08/05 00:16 E.Chudakov Shower/preshower status
  • 145466 06/08/05 00:10 cusanno y-target reconstruction, comparison with last year
  • 145465 06/08/05 00:10 cusanno y-target reconstruction, comparison with last year
  • 145463 06/08/05 00:01 cusanno CT spectrum on run 4381
  • Tuesday

  • 145461 06/07/05 23:57 sawatzky Swing shift summary
  • 145459 06/07/05 23:47 sawatzky General Tools, magnet strip charts
  • 145458 06/07/05 23:46 sawatzky BCM, SLI strip charts
  • 145455 06/07/05 23:19 cusanno Path Lenght OK
  • 145454 06/07/05 23:19 sawatzky SLI snapshot after optics 'tune'
  • 145453 06/07/05 23:15 sawatzky SLI, BCM strip chart for shift
  • 145452 06/07/05 23:12 cusanno A1 spectrum in pseudo-production
  • 145451 06/07/05 23:08 cusanno A2 in pseudo-production
  • 145449 06/07/05 23:03 sawatzky Beam is back
  • 145448 06/07/05 22:58 sawatzky MCC took beam away to run a 'quick optics plan'
  • 145446 06/07/05 22:50 sawatzky Prescale settings for run 4390+
  • 145445 06/07/05 22:43 E.Chudakov HRSL pion rejector is turned on
  • 145442 06/07/05 22:39 sawatzky SLI snapshot (50 uA production)
  • 145441 06/07/05 22:30 cusanno Y target reconstruction on waterfall
  • 145432 06/07/05 21:31 sawatzky SLI snapshot (10 uA)
  • 145431 06/07/05 21:29 sawatzky General Tools, magnet and bcm strip charts for SLI test
  • 145414 06/07/05 20:17 sawatzky SLI tests started
  • 145410 06/07/05 19:05 sawatzky General Tools, magnet strip charts
  • 145408 06/07/05 18:51 cusanno y target reconstruction on waterfall
  • 145407 06/07/05 17:57 markowitz y target shifted foil on waterfall?
  • 145406 06/07/05 17:37 Markowitz, C y-target reconstruction
  • 145405 06/07/05 17:28 R.A. LIndgre Day Shift Summary
  • 145403 06/07/05 16:26 Markowitz Run Plan for Tuesday swing, Wednesday, Thursday owl shift
  • 145402 06/07/05 15:08 LeRose Injector Problems
  • 145401 06/07/05 15:03 Holmstrom Y matrix elements
  • 145397 06/07/05 12:01 Markowitz adjusted prescale factors based on run 2200, january 2004
  • 145396 06/07/05 10:27 R. Michaels Aerogel cuts
  • 145395 06/07/05 10:17 Markowitz reboot adaql1
  • 145394 06/07/05 09:23 R.A. Lindgre Machine studies
  • 145392 06/07/05 07:57 Boeglin Shift Summary
  • 145391 06/07/05 04:56 Boeglin CHL trip
  • 145390 06/07/05 03:34 E.Chudakov HRSR shower HV change
  • 145386 06/07/05 00:45 sawatzky General Tools, magnet strip charts
  • 145383 06/07/05 00:32 sawatzky Shift summary: Monday swing
  • 145374 06/07/05 00:08 sawatzky Beam energy spread: 9x10-5
  • 145373 06/07/05 00:05 sawatzky Septum steering tests completed
  • Monday

  • 145372 06/06/05 23:09 sawatzky Q2/3 cycled on both HRSs
  • 145371 06/06/05 22:59 sawatzky Septum steering test started
  • 145370 06/06/05 22:48 sawatzky Test beam into hall
  • 145369 06/06/05 22:21 sawatzky Test beam into hall
  • 145368 06/06/05 20:52 R. Michaels BTA repair
  • 145366 06/06/05 19:43 E.Chudakov HRSR detectors - preshower - wrong database
  • 145365 06/06/05 19:32 sawatzky General Tools, Mag strip charts
  • 145364 06/06/05 18:40 sawatzky Current HRS settings
  • 145363 06/06/05 17:07 R. Feuerbach Aerogel signal NOT small
  • 145356 06/06/05 16:49 R. Michaels BTA issues
  • 145348 06/06/05 16:06 Markowitz Waterfall tacheometer calibration
  • 145347 06/06/05 15:44 Holmstrom Shift Summary - Day
  • 145346 06/06/05 15:39 Higinbotham Magnet Screen at end of 1.85 GeV run
  • 145335 06/06/05 14:39 Holmstrom dp Followup
  • 145323 06/06/05 13:19 R. Michaels T6 prelim. efficiency
  • 145316 06/06/05 12:37 Whitbeck Target Type wrong
  • 145315 06/06/05 12:35 Holmstrom dp for run 4350
  • 145312 06/06/05 12:21 Markowitz T6 efficiency and effect of changing A1 attenuation
  • 145306 06/06/05 12:06 Markowitz changed software A1, A2 gains in online database directory
  • 145304 06/06/05 11:52 Higinbotham 5 uA SLI
  • 145301 06/06/05 11:33 Holmstrom Left Arm C dp
  • 145300 06/06/05 11:32 Higinbotham General Tools for Delta=0
  • 145299 06/06/05 11:31 Higinbotham Septum for Delta=0 Runs
  • 145297 06/06/05 11:26 Holmstrom dp for both arms on Ta
  • 145289 06/06/05 09:54 Higinbotham Change to 2nd Pass
  • 145286 06/06/05 09:47 Higinbotham 90 uA SLI
  • 145283 06/06/05 09:38 Higinbotham 74 uA SLI
  • 145274 06/06/05 08:57 Higinbotham SLI for 60 uA run
  • 145273 06/06/05 08:55 Higinbotham Hall A BCM's Off from Faradey
  • 145272 06/06/05 08:52 Holmstrom Hall A tool screen
  • 145270 06/06/05 08:22 Holmstrom Q3 Not Regulating
  • 145268 06/06/05 08:12 Nilanga Shift Summary
  • 145267 06/06/05 08:05 salvatore SLI screen run 4339
  • 145262 06/06/05 07:52 salvatore SLI screen run 4337
  • 145259 06/06/05 07:36 salvatore SLI screen run 4336
  • 145256 06/06/05 07:24 salvatore SLI screen run 4335
  • 145253 06/06/05 06:47 salvatore SLI screen run 4334
  • 145250 06/06/05 05:58 salvatore SLI screen run 4333
  • 145235 06/06/05 05:04 salvatore SLI screen run 4326
  • 145228 06/06/05 04:40 Frullani SLI screen run 4323
  • 145219 06/06/05 03:23 frullani SLI screen run 4319
  • 145204 06/06/05 02:17 frullani SLI screen run 4312
  • 145191 06/06/05 01:12 frullani SLI screen run 4306
  • 145180 06/06/05 00:02 sawatzky Swing shift summary
  • Sunday

  • 145179 06/05/05 23:49 sawatzky MCC update: will be another hour or so
  • 145178 06/05/05 22:57 sawatzky General Tools, Magnet strip chart
  • 145177 06/05/05 22:49 sawatzky MCC: lost power to a service building
  • 145171 06/05/05 22:18 markowitz lhrs is back
  • 145170 06/05/05 22:04 Markowitz TECH left hrs ioc crash
  • 145169 06/05/05 21:11 R. Michaels pedrun fixed
  • 145168 06/05/05 20:33 sawatzky Right septum quenched
  • 145167 06/05/05 20:28 sawatzky CODA Pedestal configuration not working
  • 145166 06/05/05 20:15 sawatzky Changes to CODA 'epicsRunEnd' script
  • 145165 06/05/05 19:17 Markowitz Sieve slits in, changed T6 A1 attenuation
  • 145163 06/05/05 17:35 sawatzky 1.795 GeV/c Septum strip chart
  • 145160 06/05/05 17:22 sawatzky 1.795 GeV/c General tools, magnet strip chart
  • 145157 06/05/05 16:28 Markowitz day shift summary
  • 145139 06/05/05 12:01 Armando Acha dp/p =3E-5 Sigma L HRS
  • 145137 06/05/05 11:54 markowitz tools, strip charts at 1.776 GeV/c (-4%)
  • 145136 06/05/05 11:53 markowitz septum screens for 1.776 GeV
  • 145135 06/05/05 11:31 Armando Acha dp/p =3E-5 sigma R HRS
  • 145131 06/05/05 10:49 Armando Acha VDC tracking reference HALog 145105
  • 145128 06/05/05 10:35 LeRose Septum Setting Revised
  • 145126 06/05/05 10:07 markowitz septum screens 1.813 GeV -2%
  • 145125 06/05/05 10:06 markowitz tools, strip charts for 1.813 GeV -2%
  • 145124 06/05/05 09:50 salvatore owl shift summary
  • 145121 06/05/05 09:14 markowitz main tools screen for 1.850 GeV
  • 145117 06/05/05 09:08 Markowitz added iocha49 to epics Run Start/End
  • 145116 06/05/05 09:01 Markowitz run start and stop changes to epics logger
  • 145115 06/05/05 08:55 Markowitz run start and stop changes to epics logger
  • 145113 06/05/05 08:48 Markowitz magnet strip charts 1.850 GeV/c
  • 145112 06/05/05 08:43 markowitz septum screens for 1.850
  • 145108 06/05/05 07:51 salvatore screens +2.0% sieve out
  • 145105 06/05/05 07:10 cusanno VDC tracking with no sieve
  • 145102 06/05/05 06:51 salvatore screens +4.0% sieve out
  • 145087 06/05/05 05:06 salvatore screens -4.0%
  • 145082 06/05/05 04:18 salvatore screens -2.0%
  • 145077 06/05/05 03:33 salvatore screens 0.0%
  • 145072 06/05/05 02:41 salvatore screens +2.0%
  • 145070 06/05/05 02:36 salvatore wrong commentfor 145059 entry
  • 145068 06/05/05 02:02 sawatzky Shift Workers: Update run-start text each run
  • 145065 06/05/05 01:53 salvatore wrong information for FSD masks
  • 145060 06/05/05 01:34 salvatore delta scan 4.0%
  • 145058 06/05/05 00:07 sawatzky Runs 4250--4253 are JUNK
  • 145057 06/05/05 00:04 sawatzky Swing shift summary
  • Saturday

  • 145054 06/04/05 23:56 LeRose Momentum Changes
  • 145052 06/04/05 23:41 LeRose Septum Magnet Current
  • 145050 06/04/05 23:33 adaq halla general tools
  • 145048 06/04/05 23:31 LeRose MCC Plans
  • 145043 06/04/05 22:38 adaq general halla tools
  • 145042 06/04/05 22:28 Ramesh Some changes need to be made in some scripts
  • 145041 06/04/05 22:27 sawatzky MCC's owl shift beam plans
  • 145037 06/04/05 21:33 Brad/Ramesh Effect of Left septum current adjustment
  • 145029 06/04/05 20:36 Ramesh/Rob +ve and -ve polarity of right Q1
  • 145027 06/04/05 20:33 sawatzky General tools, Magnet stripchart
  • 145021 06/04/05 19:50 Ramesh/Brad right septum on and off
  • 145013 06/04/05 18:58 adaq hall a general tool
  • 145006 06/04/05 18:17 sawatzky Add waterfall target to CODA run-start dialog
  • 144996 06/04/05 17:32 LeRose Septum Setting
  • 144994 06/04/05 17:28 Ramesh Sieve picture
  • 144988 06/04/05 16:38 folts tech
  • 144987 06/04/05 16:12 E.Chudakov Day shift summary
  • 144986 06/04/05 16:05 E.Chudakov Both arms quenched
  • 144983 06/04/05 15:41 LeRose Right Septum Quench
  • 144981 06/04/05 15:35 E.Chudakov Septa OK now!
  • 144970 06/04/05 15:04 E.Chudakov Beam parameters
  • 144965 06/04/05 14:48 E.Chudakov Beam is back
  • 144962 06/04/05 14:33 E.Chudakov MCC reported problems restoring our beam
  • 144961 06/04/05 13:16 LeRose Septum Magnet Setting
  • 144960 06/04/05 13:09 E.Chudakov End of control access
  • 144958 06/04/05 11:27 E.Chudakov Septa polarity seems wrong
  • 144952 06/04/05 08:58 E.Chudakov Magnet strip chart, main screen
  • 144949 06/04/05 08:44 Markowitz spot
  • 144946 06/04/05 08:29 markowitz arc scan
  • 144945 06/04/05 08:11 Markowitz owl shift summary
  • 144943 06/04/05 07:21 markowitz harp scan
  • 144942 06/04/05 07:01 Markowitz 2 pass beam being setup for hall A
  • 144941 06/04/05 01:27
  • Friday

  • 144938 06/03/05 22:41
  • 144929 06/03/05 18:09 LeRose VDC's on
  • 144928 06/03/05 18:07 LeRose Status/Plan
  • 144923 06/03/05 15:53 R. Michaels cables to reconnect
  • 144922 06/03/05 15:50 adaq a test
  • 144917 06/03/05 15:27 LeRose It works now!
  • 144916 06/03/05 15:20 LeRose test entry
  • 144915 06/03/05 15:13 LeRose Test Entry
  • 144914 06/03/05 15:05 LeRose Right Arm Field Change
  • 144912 06/03/05 13:28 LeRose Left VDC's OFF
  • Thursday

  • 144910 06/02/05 23:57 Ramesh Everything ready for elastic-kinematics except the beam
  • 144909 06/02/05 19:42 Ramesh Hall A general tools, left hrs magnets are great!
  • 144908 06/02/05 19:40 Ramesh holog entries made from adaql8 did not go to the web
  • 144907 06/02/05 18:22 LeRose Status
  • 144904 06/02/05 17:11 Spiegel Hall A Tech's
  • 144898 06/02/05 06:54 Markowitz Shift/ beam steering summary
  • 144897 06/02/05 05:01 Markowitz both septum, both Q1
  • 144896 06/02/05 04:51 Markowitz Right Septum ,Q1and Left Septum
  • 144895 06/02/05 04:34 Markowitz right Q1 and septum at 2 GeV
  • 144894 06/02/05 04:26 Markowitz right septum at 185 amps
  • 144893 06/02/05 04:06 Markowitz start of steering tests
  • 144892 06/02/05 04:01 Markowitz target move to empty
  • 144891 06/02/05 03:39 Markowitz Target motion (12C thin)
  • Wednesday

  • 144887 06/01/05 17:41 R. Michaels TS1 needs to be power cycled